# zkPHIRE: A Programmable Accelerator for ZKPs over HIgh-degRee, Expressive Gates

### v1

@@ III-A Challenges
At its core, SumCheck offers many degrees of parallelism. Different MLE tables can compute their extensions in parallel, extensions for different MLE pairs within a table can be computed in parallel, and the products across different extensions can be computed in parallel. Given a SumCheck polynomial, an optimal circuit can be constructed to maximize reuse and throughput, as done in zkSpeed [12].
However, building a generalizable SumCheck unit poses several challenges. For one, the unit must support a diverse set of composite polynomial structures used to express different types of constraints or implement different kinds of operations. The unit must support an arbitrary number of terms and an arbitrary degree. This means that the optimal resource allocation (e.g., number of modular multipliers) for one composite polynomial may be suboptimal or wasteful for another composite polynomial. Additionally, many composite polynomials have individual polynomials that repeat. For example, in $f=a\cdot b\cdot e+c\cdot e+e\cdot g$, the $e$ polynomial appears three times. While we could parallelize all computations for just the first term ($a\cdot b\cdot e$) before proceeding to the next term, this would leave out a clear opportunity for data reuse. Indeed, zkSpeed’s design captures immediate reuse, but is tailored to HyperPlonk’s vanilla gate polynomials. Their design philosophy overprovisions resources for simpler polynomials, and is unable to support complex polynomials.
Figure 1: The SumCheck dataflow of size-$2^{3}$ polynomial $f(X)=abce$. For a degree-4 polynomial $f$, every MLE evaluation pair is used to extend to five evaluations. Extended evaluations are multiplied, summed, and hashed to get a challenge that is used to update MLE tables for the next SumCheck round.
@@ III-B Datapath
Our programmable SumCheck unit seeks to balance the tradeoff between resource allocation and capturing immediate data reuse. Figure 4 shows the datapath for our programmable SumCheck unit. It consists of several processing elements (PEs), with MLE Update units, Extension Engines (EEs) and Product Lanes. Instead of streaming all individual polynomials’ MLE entries simultaneously, we proceed one term at a time, fetching a tile of data from each MLE table and storing in local scratchpad buffers. We allocate 16 scratchpad buffers, more than sufficient to accommodate polynomial structures we see in current ZKP systems. The scratchpad buffers are banked to match the number of PEs for parallel accesses. This ensures that individual polynomials reused in subsequent terms need not be fetched from off-chip memory for a given tile.
For a given term, the polynomials are read into the Update units and EEs, where their extensions are generated by a series of modular adders and subtractors. Specifically, in round 1 of SumCheck, two values are read at a time per MLE and directly fed to generate extensions. In all subsequent rounds, four values are read per MLE, because we pipeline the updates into the extension computation. Updating with four inputs ensures that two values are generated and fed to the EEs to match the throughput of round 1. These updated values are also written to FIFOs that buffer writebacks to off-chip, forming the (halved) tables for the next SumCheck round. In later rounds of SumCheck, when MLE tables can fit fully on-chip, writebacks to off-chip are eliminated and updated values are directly routed to the scratchpad banks.
After the extensions for each MLE are generated, they must be packed into their respective groups, e.g., all extensions at 1 are routed together. Each group of extensions is then fed into a Product Lane (PL). Each PL is equipped with E1E-1 fully pipelined modular multipliers, where $E$ is the number of EEs. Each PL’s output is accumulated into a register corresponding to the extension, e.g., all products at $X_{i}=2$ are always routed to register $2$. We allocate 32 registers (up to degree 31); higher degrees are supported by storing in scratchpads.
@@ III-C Mapping Polynomials to EEs
We devise a scheduler that determines which MLE tables are fed to EEs and PLs using a graph-based decomposition (see Figure 2). In this example, we assume there are 3 EEs, so only 3 MLEs can be handled on-chip at a time. The polynomial is adding a degree-6 term and a degree-3 term. Since there are more MLEs than EEs, we must store the partial products for all extensions; this is done using a dedicated Tmp MLE buffer.
A natural way to schedule the MLEs is using the tree-based approach on the left in Figure 2. However, this approach would require storing two distinct intermediate products that would need to be combined. As the polynomial degree grows, this tree-based approach would require more and more intermediate buffers. Instead, we can use the accumulation based schedule on the right. This scheme only requires 1 Tmp MLE buffer that accumulates extension products within the same term. Notice that this scheme uses the same number of steps as the balanced tree while minimizing temporary storage.
Figure 2: Graph decomposition used to schedule computation on our SumCheck unit. The method on the right minimizes on-chip temporary MLE buffer size. Numbers inside each node indicate the step in the schedule. When processing term hknhkn, $h$ is already prefetched during the prior step.
In this way, the scheduler traverses the polynomial term-by-term, deciding which MLEs are handled by the PEs in each step. These steps are programmed into a controller on-chip. In each step $j$, the PEs read the MLEs identified for step $j$. The controller simultaneously prefetches a tile of data for each MLE in step $j+1$, but only if the MLEs have not already been fetched on-chip. For example, the $h$ MLE would be prefetched during step $2$ (on the right), and the only $k$ and $n$ MLEs would be prefetched during step $3$. Another advantage of this scheme is that prefetch bandwidth is more balanced over steps.
@@ III-D Mapping Extensions to PLs
Since our SumCheck unit must handle any number of extensions, we may run into scenarios where there are $K$ extensions but only $P<K$ lanes. In this case, we have an initiation interval $II=K/P$ and use delay buffers to queue unmatched extensions. In Figure 4, with $K=5$ and $P=3$, the first cycle maps extensions 022 to PLs, while extensions 3344 are buffered for the next cycle. Scheduling then proceeds in a pipelined fashion, with buffered and incoming extensions interleaved across PLs. While $K$ depends on polynomial degree, the schedule for a given $(K,P)$ is static and programmed into the controller at the start of each SumCheck.
III-E Built-in support for ZeroChecks
Many ZKP protocols use SumCheck to verify circuit correctness by checking that a polynomial $f(X)$ evaluates to zero across all inputs. However, simply verifying $\sum_{X}f(X)=0$ is insufficient. Individual gates could evaluate to nonzero values (meaning they were incorrectly executed), but cancel each other out, so the total sum would still be zero. To prevent this, protocols like [50, 61, 7] use ZeroCheck, which multiplies $f(X)$ with a randomized auxiliary polynomial $f_{r}(X)$ and checks that $\sum_{X}(f(X)\cdot f_{r}(X))=0$. $f_{r}(X)$ is constructed on-the-fly from a vector of random challenges generated by a SHA module (referred to as Build MLE in zkSpeed), and zkPHIRE integrates this auxiliary polynomial construction directly into the SumCheck datapath. During the first round, one product lane is dedicated to computing $f_{r}(X)$, with its outputs routed to scratchpad banks and the EEs in parallel with the fetched MLE elements. This reduces the number of lanes for product computations by one in that round but avoids separate $O(N)$ precomputation and memory overhead, reusing existing PL multipliers to mask the cost.
@@ III-E Built-in support for ZeroChecks
We include the programmable SumCheck unit as part of a larger accelerator framework for the HyperPlonk protocol. We describe the steps of the HyperPlonk protocol and its mapping onto zkPHIRE units. HyperPlonk is based on the BLS12-381 elliptic curve, which uses 255-bit scalars (datatypes used by MLEs) and 381-bit elliptic curve points (used for point addition in MSMs). We explore the use of both arbitrary and fixed prime modular multipliers (as done in [49]); the latter saves us roughly 50% on area and improves our computational density by roughly $2\times$. Figure 4 shows the overall architecture.
@@ IV-A HyperPlonk Protocol Overview
HyperPlonk has a very complex dataflow. Here, we describe the high-level protocol and what hardware modules each step requires. The 5 main steps are Witness Commitments, Gate Identity Check, Wire Identity Check, Batch Evaluations, and Polynomial Opening. In Witness Commitments, the prover commits to their private input values (witnesses) using multi-scalar multiplications (MSMs) so that they cannot claim other witness values later. We run this step with the MSM module. In Gate Identity Checks, the prover shows that all gates in the circuit were computed correctly via ZeroChecks. We run this step with our SumCheck unit and a Multifunction Forest unit. In Wire Identity Checks, the prover proves that the wiring between gates is consistent with the circuit’s intended connections. This is done by constructing wiring permutation polynomials and running a PermutationCheck (henceforth called a PermCheck) and then committing to those polynomials. We run this step with a PermCheck Generator unit, the SumCheck and Multifunction Forest units, and the MSM unit. In Batch Evaluations, the prover evaluates all committed polynomials at various points provided by the verifier. We run this step with the Multifunction Forest unit. In the Polynomial Opening, the prover reveals the values of multiple committed polynomials at specific points and proves that it is consistent with the original commitment. The prover then runs a SumCheck to combine these polynomial evaluation claims into one (we call this OpenCheck as per zkSpeed). The combined polynomial commitment is then opened using the MSM unit. Overall this step uses an MLE Combine unit, the SumCheck and Multifunction Forest units, and the MSM unit.
Masking ZeroCheck: In zkSpeed and the HyperPlonk baseline, Gate Identity and Wire Identity steps run serially, despite verifying independent circuit properties. zkPHIRE includes support to overlap the Gate Identity’s ZeroCheck with the MSMs during Wiring Identity. Since MSMs dominate runtime and have low bandwidth pressure due to high data reuse [13, 12], this scheduling effectively masks ZeroCheck latency.
@@ IV-B Accelerator Units
IV-B1 Sparsity Handling for SumCheck
@@
zkSpeed opts to use large on-chip scratchpads to store these MLEs, because they are reused throughout steps of the HyperPlonk protocol. However, as the problem size scales, the area consumed by the scratchpad grows, with 10s of MBs dedicated to address translation buffers alone, and a total on-chip storage of nearly 300 MB. Furthermore, the MLEs are only useful for round 1 of SumChecks. For subsequent rounds, the MLEs are updated and written to off-chip memory, so the global scratchpad is idle for roughly 50% of the SumCheck runtime. zkPHIRE opts to use smaller scratchpad buffers to store tiles of each MLE. While this results in more accesses to off-chip memory and we incur more fill/drain latency penalties, it allows for more area to be dedicated core computation structures within SumCheck and other accelerator modules. The improved speedups offset the latency hits incurred by eliminating large scratchpads. To eliminate address translation overheads, zkPHIRE fetches lightweight per-tile offset buffers for witness MLEs. The SumCheck controller decodes 255-bit elements using these offsets, while binary MLEs are stored directly without translation. This scheme incurs minimal bandwidth overhead while enabling more compact storage.
@@ IV-B2 Multifunction Forest
In HyperPlonk, there are several sub-steps that can be mapped to a binary-tree computation, including MLE evaluation, building MLEs, computing the product MLE ($\pi$), and multiplication tree. We adopt the same base tree unit architecture as zkSpeed’s MTU [38]. However, unlike zkSpeed – which employs a single tree unit for all sub-step computations while maintaining a separate, dedicated SumCheck module – zkPHIRE introduces a more unified and efficient approach. In zkPHIRE, we implement the product lanes of the programmable SumCheck unit using a multifunction forest. This allows modular multipliers initially provisioned just for SumCheck to be shared across other protocol steps. The shared multifunction forest optimization helps zkPHIRE achieve the same latency as zkSpeed at iso-workload-length with 15.2% fewer modular multipliers.
@@ IV-B3 MSM
We use the same MSM architecture as zkSpeed, since we target the same protocol, HyperPlonk. In HyperPlonk, the number of MSMs only depends on the number of witnesses. The underlying hardware is unchanged, we just compute more MSMs using it. Jellyfish gates have five witnesses, so we run five Sparse MSMs and three Dense MSMs.
@@ IV-B4 MLE Combine
IV-B5 PermCheck Generator
Figure 5: The PermCheck Generator Module.
PermCheck requires constructing Numerator (N), Denominator (D), Fraction ($\phi$), and Product ($\pi$) MLEs [12]. Since $\phi$ is computed on-the-fly from N and D, we design a pipelined unit that generates all three in parallel, producing one element per cycle after warmup (Figure 5). The Multifunction Forest is used to construct $\pi$ from $\phi$, as in zkSpeed. To accommodate both vanilla and custom gates, we allocate 5 PEs (one per witness), with the ability to support more than 5 through overlapped scheduling and cyclic reuse of PEs. Intermediate $N_{1\dots k}$ and $D_{1\dots k}$ MLEs are written to HBM and combined via modular multiplications to produce the final N and D MLEs, followed by modular inversion of D.

### v2

@@ III-A Challenges
At its core, SumCheck offers several degrees of parallelism. Different MLE tables can compute their extensions in parallel, extensions for different MLE pairs within a table can be computed in parallel, and the products across different extensions can be computed in parallel. Given a SumCheck polynomial, an optimal circuit can be constructed to maximize reuse and throughput, as done in zkSpeed [9].
In contrast, a generalizable SumCheck unit must support a diverse set of composite polynomial structures used to express different types of constraints or implement different kinds of operations. The unit must support an arbitrary number of terms and an arbitrary degree. This means that the optimal resource allocation (e.g., number of modular multipliers) for one composite polynomial may be suboptimal or wasteful for another one. Additionally, many composite polynomials have individual polynomials that repeat. For example, in $f=a\cdot b\cdot e+c\cdot e+e\cdot g$, the $e$ polynomial appears three times. While we could parallelize all computations for the first term ($a\cdot b\cdot e$) before proceeding to the next term, this would leave out a clear opportunity for data reuse. Indeed, zkSpeed’s design captures immediate reuse, but is tailored to HyperPlonk’s Vanilla gate polynomials. zkSpeed overprovisions resources for simple polynomials, and cannot support complex composite polynomials.
Figure 1: The SumCheck dataflow of size-$2^{3}$ polynomial $f(X)=abce$. For a degree-4 polynomial $f$, every MLE evaluation pair is extended to five evaluations. Extended evaluations are multiplied, summed, and hashed to get a challenge that is used to update MLE tables for the next SumCheck round.
@@ III-B Datapath
Our programmable SumCheck unit seeks to balance the trade-off between resource allocation and capturing immediate data reuse. Figure 4 shows the datapath for our programmable SumCheck unit. It consists of several processing elements (PEs), with MLE Update units, Extension Engines (EEs) and Product Lanes (PLs). Instead of streaming all individual polynomials’ MLE entries simultaneously, we proceed one term at a time, fetching a tile of data from each MLE table and storing in local scratchpad buffers. We allocate 16 scratchpad buffers, more than sufficient to accommodate polynomial structures we see in current ZKP systems. The scratchpad buffers are banked to match the number of PEs for parallel accesses. This ensures that individual polynomials reused in subsequent terms need not be fetched from off-chip memory for a given tile.
For a given term, the polynomials are read into the Update units and EEs, where their extensions are generated by a series of modular adders and subtractors. Specifically, in round 1 of SumCheck, each PE reads two values at a time from MLE tables and directly feeds them to the EEs to generate extensions. In all subsequent rounds, four values are read per MLE, because we pipeline the MLE updates directly into the extension computation. Updating with four inputs ensures that two values are generated and fed to the EEs to match the same throughput of round 1. These updated values are also written to FIFOs that buffer writebacks to off-chip, forming the (halved) tables for the next SumCheck round. In later rounds of SumCheck where MLE tables can fit fully on-chip, writebacks to off-chip are eliminated and updated values are directly routed to the scratchpad banks.
After each MLE’s extensions are generated, they are packed into their respective groups, e.g., the extensions for all MLEs at $X_{i}=2$ are routed together. Each group of extensions is then fed into a Product Lane (PL). Each PL is equipped with E1E-1 fully pipelined modular multipliers, where $E$ is the number of EEs. Each PL’s output is accumulated into a register corresponding to the extension, e.g., all products at $X_{i}=2$ are routed to register $2$. We allocate 32 registers (up to degree 31); higher degrees are supported by storing in scratchpads.
@@ III-C Mapping Polynomials to EEs
We devise a scheduler that determines which MLE tables are fed to EEs and PLs using a graph-based decomposition (see Figure 2). In this example, we assume there are 3 EEs, so only 3 MLEs can be handled by a SumCheck PE at a time. The polynomial is adding a degree-6 term and a degree-3 term. Since there are more MLEs than EEs, we must store the intermediate products for all extensions, not just the initial extensions at $X_{i}=0,1$ that are provided from the original MLE tables. This is done using a dedicated Tmp MLE buffer.
A natural way to schedule the MLEs is using the balanced tree-based approach on the left in Figure 2. However, this would require storing two distinct intermediate products that would need to be combined. As the polynomial degree grows, this tree-based approach would require more and more intermediate buffers. Instead, we can use the accumulation-based schedule on the right. This scheme only requires 1 Tmp MLE buffer that accumulates extension products within the same term. Notice that this scheme uses the same number of steps as the balanced tree while minimizing temporary storage.
Figure 2: Graph decomposition used to schedule computation on our SumCheck unit. The method on the right minimizes the number of temporary MLE buffers. Numbers inside each node indicate the step in the schedule. When processing term hknhkn, $h$ is already prefetched during the prior step.
This way, the scheduler traverses the polynomial term-by-term, deciding which MLEs are handled by the PEs in each step. In each step $j$, the PEs read the MLEs assigned to step $j$, while a tile of data for each MLE needed in step $j+1$ is prefetched if it is not already brought on-chip. For example, the $h$ MLE is prefetched during step $2$ (on the right), while only the $k$ and $n$ MLEs are prefetched during step $3$. For higher-degree polynomials, this scheme balances prefetch bandwidth across steps rather than front-loading it in the initial steps, as would occur with a binary-tree schedule.
@@ III-D Mapping Extensions to PLs
Our SumCheck unit must handle any number of extensions, and we may run into scenarios where there are $K$ extensions but only $P<K$ lanes. In this case, we have an initiation interval $II=K/P$ and use delay buffers to queue extra extensions. In Figure 4, with $K=5$ and $P=3$, the first cycle maps extensions 022 to PLs, while extensions 3344 are buffered for the next cycle. Scheduling proceeds in a pipelined fashion, with buffered and incoming extensions interleaved across PLs.
III-E Programming the SumCheck unit
The SumCheck unit is programmed via an automated scheduler given the target polynomial and hardware configuration. Using the approach shown in Figure 2, the scheduler generates a list of computational steps, including MLE-to-EE mappings, prefetch ordering, and schedules for specific $(K,P)$ settings. These steps are annotated with signals for control registers (e.g., MLE bank selection, arbitration, bypassing update), address offsets, and FSM configuration. They are then loaded into on-chip controllers as instructions to guide SumCheck execution.
III-F Built-in support for ZeroChecks
Many ZKP protocols use SumCheck to verify circuit correctness by checking that a polynomial $f(X)$ evaluates to zero across all inputs. However, simply verifying $\sum_{X}f(X)=0$ is insufficient. Individual gates could evaluate to nonzero values (meaning they were incorrectly executed), but cancel each other out, so the total sum would still be zero. To prevent this, protocols like [55, 67, 4] use ZeroCheck, which multiplies $f(X)$ with a randomized auxiliary polynomial $f_{r}(X)$  [^1] In the ZKP literature [55, 6, 3], $f_{r}(x)$ is written as $eq(x,r)$ and checks that $\sum_{X}(f(X)\cdot f_{r}(X))=0$. $f_{r}(X)$ is constructed on-the-fly from a vector of random challenges generated by a SHA module. The construction of an $f_{r}(X)$ polynomial is referred to as the Build MLE kernel in zkSpeed, and we integrate this auxiliary polynomial construction directly into the SumCheck datapath. During the first SumCheck round, one product lane is dedicated to computing $f_{r}(X)$. Its outputs are routed to a dedicated scratchpad bank and one EE, while the remaining EEs fetch MLE elements from the other scratchpad banks. One less lane is available for product computations for the first SumCheck round, but this avoids separate $O(N)$ precomputation of $f_{r}$ and the memory overhead of writing and reading back values. Subsequent SumCheck rounds treat $f_{r}$ as any other MLE fetched from and updated to off-chip.
@@ III-F Built-in support for ZeroChecks
We include the programmable SumCheck unit as part of zkPHIRE, a larger accelerator framework for the HyperPlonk protocol. Figure 4 shows the overall architecture.
@@ IV-A HyperPlonk Protocol Overview
HyperPlonk has a very complex dataflow as outlined in [9]. Here, we describe the high-level protocol and which hardware modules each step requires. The 5 main steps are Witness Commitments, Gate Identity Check, Wire Identity Check, Batch Evaluations, and Polynomial Opening.
In Witness Commitments, the prover commits to their private input values (witnesses) using multi-scalar multiplications (MSMs) so that they cannot claim other witness values later. We run this step with the MSM unit. In Gate Identity Checks, the prover shows that all gates in the circuit were computed correctly via ZeroChecks. We run this step with our SumCheck unit and a Multifunction Forest unit. In Wire Identity Checks, the prover proves that the wiring between gates is consistent with the circuit’s intended connections by enforcing a permutation argument over the wire values. This is done by constructing wiring permutation polynomials and forming a quotient and grand product argument that encodes the permutation constraint. These polynomials are committed to, and then correctness of the grand product is verified via a PermutationCheck (PermCheck). We run this step with a Permutation Quotient Generator unit, the SumCheck and Multifunction Forest units, and the MSM unit.
In Batch Evaluations, the prover evaluates all committed polynomials at various points provided by the verifier. We run this step with the Multifunction Forest unit. In the Polynomial Opening, the prover reveals the values of multiple committed polynomials at specific points and proves that it is consistent with the original commitment. The prover then runs a SumCheck to combine these polynomial evaluation claims into one (we call this OpenCheck as per zkSpeed). The combined polynomial commitment is then opened using the MSM unit. Overall this step uses an MLE Combine unit, the SumCheck and Multifunction Forest units, and the MSM unit.
Masking ZeroCheck: In zkSpeed and the HyperPlonk baseline, Gate Identity and Wire Identity steps run serially, despite verifying independent circuit properties. zkPHIRE includes support to overlap the Gate Identity’s ZeroCheck with the MSMs during Wiring Identity. MSMs dominate runtime and have low bandwidth pressure due to high data reuse [10, 9], so this scheduling effectively masks ZeroCheck latency.
@@ IV-B Accelerator Units
IV-B1 SumCheck Unit with Sparsity Handling
@@
zkSpeed opts to use large on-chip scratchpads to store these MLEs, because they are reused throughout steps of the HyperPlonk protocol. However, as the problem size scales, the area consumed by the scratchpad grows, with 10s of MBs dedicated to address translation buffers alone, and a total on-chip storage of nearly 300 MB. Furthermore, the MLEs are only useful for round 1 of SumChecks. For subsequent rounds, the MLEs are updated and written to off-chip memory, so the global scratchpad is idle for roughly 50% of the SumCheck runtime. zkPHIRE opts to use smaller scratchpad buffers to store tiles of each MLE. While this results in more access to off-chip memory and incurs more fill/drain latency penalties, it allows for more area to be dedicated to core computational structures. The improved speedups offset the latency hits incurred by using smaller scratchpads.
zkSpeed’s scratchpads compress MLEs but require address translation units to decode when MLE entries are sparse or dense. zkPHIRE eliminates the address translation overhead by fetching per-tile offset buffers for witness MLEs. For a tile of data fetched from off-chip, the SumCheck controller uses the offset buffer to determine where 255-bit elements begin in the bitstream; all other bits are interpreted as $0/1$ MLE entries and written into the corresponding addresses in the scratchpads. The enable MLEs are stored as-is without using address translation. These optimizations incur minimal bandwidth costs while enabling more compact storage.
@@ IV-B2 Multifunction Forest
In HyperPlonk, several sub-steps naturally map to binary-tree computations, such as MLE evaluation, computing the product MLE ($\pi$), and building MLEs [40]. We adopt the same base tree architecture as zkSpeed’s MTU [40, 9], but differ in how it is integrated. Whereas zkSpeed uses a single tree unit alongside a dedicated SumCheck module, zkPHIRE instantiates multiple tree units as a Multifunction Forest. These units are reused to implement product-lane computations, removing the need for the extra modular multipliers in zkSpeed’s SumCheck module while keeping the pipeline fully utilized. This shared design allows the same multipliers to serve both SumCheck products and other tree-based operations, improving utilization and reducing area. As a result, zkPHIRE achieves the same latency as zkSpeed for the same workload with 15% fewer multipliers.
@@ IV-B3 MSM
We use the same MSM architecture as zkSpeed, since we target the same protocol, HyperPlonk. In HyperPlonk, the number of MSMs only depends on the number of witnesses. The underlying hardware is unchanged, we just compute more MSMs using it. Jellyfish gates have 5 witnesses, so we run 5 Sparse MSMs and 3 Dense MSMs.
@@ IV-B4 MLE Combine
IV-B5 Permutation Quotient Generator
Figure 5: The Permutation Quotient Generator Module.
PermCheck requires constructing Numerator ($N$), Denominator ($D$), Fraction ($\phi$), and Product ($\pi$) MLEs [9]. Since $\phi$ is computed on-the-fly from N and D, we design a pipelined unit that generates all three simultaneously, producing one element per cycle after warmup (Figure 5). The Multifunction Forest is used to construct $\pi$ from $\phi$, as in zkSpeed. To accommodate both Vanilla and custom gates, we allocate 5 PEs (one per witness), with the ability to support more than 5 through overlapped scheduling and cyclic reuse of PEs. Intermediate $N_{1\dots k}$ and $D_{1\dots k}$ MLEs are written to HBM and combined with modular multiplications to produce the final $N$ and $D$ MLEs, followed by modular inversion of $D$.